OpenStudy (anonymous):

cosx=5/13 and sin y=3/5 what is the exact value of cos y?

OpenStudy (anonymous):

4/5

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

sin y = opposite side / hypotenuse = 3/5 cos y = adjacent side/ hypotenuse from pythagoras theorem, adjacent side = sqrt(hypotenuse^2-opposite side^2) = sqrt(5^2-3^2) = 4 so cos y = 4/5